OpenStudy (anonymous):

What is the distance from (-2, -4) to (1, 0)?

OpenStudy (lgbasallote):

use distance formula \[\sqrt{x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\implies \sqrt{(1 - (-2) )^2 + (0 - (-4) )^2}\] \[\implies \sqrt{(1+2)^2 + (0 + 4)^2}\] got it?

OpenStudy (anonymous):

is that the awnser

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

you have to solve it further

OpenStudy (anonymous):

i have no idea..3 units 4 units 5 units 6 units

OpenStudy (anonymous):

(1+2) squared = 9 and (0 + 4) squared = 16. And now you add the two answers and you get 25. the last step you take the square root of 25. please tell me you know how to get that

OpenStudy (anonymous):

no..